5 years ago, Jay was seven times older than Mary. In five years, Mary will be half as old as Jay (or Jay will be twice as old as Mary). How old is each one?
Answer by stanbon(75887)

5 years ago, Jay was seven times older than Mary. In five years, Mary will be half as old as Jay (or Jay will be twice as old as Mary). How old is each one?
------------------------
Equations:
(j-5) = 7(m-5)
------
(m+5)= (1/2)j
-----------------------
Simplify each equation:
j - 7m = -30
j - 2m = 10
----------------------
Substitute and solve for "m":
2m+10 = 7m-30
5m = 40
m = 8 (Mary's age now)
----
Since j = 2m+10, j = 26 (Jay's age now)
=========================================
Cheers,
Stan H.
=========================================